41.Suppose that )-for o < lowing probabilities: (a) P(l < X) (b) P <X<2.5) (c)P(X= 3) (d) P(X< 4) (e) P3 s x) Determine the fol-
Answer #1

4.1)

this is exponential distribution

P(X > x) =e^(- lambda *x)

here lambda = 1

a)

P(X >1) = e^(-1)

b)

P(1 <X< 2.5) = e^(-1) - e^(-2.5)

c)

P(X = 3) = 0

d)

P(X < 4) = 1- e^(-4)

e)

P(X >= 3) = e^(-3)

Please rate